# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v4

@@ 2. $(2)$
    or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$.
@@
Since $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, we have $u(\mathbf{x}^{\prime})=u(\mathbf{x})_{[n]\setminus(D_{1}\cup D_{2})}$. It follows that $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$. Suppose that $U(\mathbf{x}^{\prime})_{1}$ is obtained from $U(\mathbf{x})_{1}$ by deleting $U(\mathbf{x})_{1,k_{1}}$ and $U(\mathbf{x})_{1,k_{2}}$. Recall that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. So $U(\mathbf{x}^{\prime})_{1}$ is obtained from $u(\mathbf{x})_{I_{n,b}}$ by deleting $u(\mathbf{x})_{1+(k_{1}-1)b}$ and $u(\mathbf{x})_{1+(k_{2}-1)b}$.
Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.2, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$. In addition, by Lemma IV.3, we can
@@
  • or find an interval $[c,d]\subseteq[n/b]$ of length at most $3d\log(n/b)$ such that $k_{1},k_{2}\in[c,d]$.
By II.2, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. For the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$. Now the proof is completed. ∎
From now on, we assume that $\mathbf{x}\in\Sigma_{q}^{n}$ and $A(\mathbf{x})_{1}$ is $d$-regular, where $d$ will be specified later. Furthermore, suppose $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$. Then by II.2, we have $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$. Since $U(\mathbf{x})_{1}$ is regular, Lemma IV.2 ensures that when given $\eta\left(U(\mathbf{x})_{1}\right)$, one can uniquely recover $U(\mathbf{x})_{1}$ from $U\left(\mathbf{x}^{\prime}\right)_{1}$. Then by Lemma IV.3, we can know the approximate positions of the two deletions in $U(\mathbf{x})_{1}$. According to Lemma IV.5, we can approximately determine the positions of the two bursts in $\mathbf{x}$. Based on the two cases in Lemma IV.5, we split our discussion into Section IV-B and Section IV-C. Our codes for correcting two $b$-burst-deletions are presented in Theorem IV.1, whose proof is divided into Lemmas IV.6 and IV.7.
@@
Let $m>b$. Suppose $\mathbf{u}\in\Sigma_{q}^{m}$ and $\mathbf{v}\in\mathcal{B}_{1}^{b}(\mathbf{u})$. Then for any $b\leq i\leq m-b$, we have $\mathbf{v}_{[1,i-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[1,i]}\right)$ and $\mathbf{v}_{[i+1,m-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[i+1,m]}\right)$.
@@
In particular, $\mathbf{x}_{I_{j}^{\prime}}$ is know to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, to decode $\mathbf{x}$ from $\mathbf{x}^{\prime}$, it is sufficient to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
@@
Since $j_{2}-j_{1}\geq 2$, deletions in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ do not affect $\mathbf{x}_{I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}}$. As a result, (13) implies $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}\right)$ and $\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}}\right)$. Applying IV.1 to $\mathbf{u}=\mathbf{x}_{I_{j_{1}-1}\cup I_{j_{1}}}$, $\mathbf{v}=\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$, $m=\left(p_{j_{1}}-p_{j_{1}-2}\right)b$ and $i=\left(p_{j_{1}-1}-p_{j_{1}-2}\right)b$, we conclude that
@@
Recall that the range of $j_{1}$ is $\left[1,r(U(\mathbf{x})_{1})\right]$. To make (14) valid when $j_{1}=1$, we let $\mathbf{x}_{I_{0}^{\prime}}=0^{b}$ (the sequence consisting of $b$ symbols 00) and $\mathbf{x}^{\prime}_{\left[p_{-2}b+1,-b\right]}$ be an empty sequence.
By (14) and (15), to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$, we need to apply to each $\mathbf{x}_{I_{j}^{\prime}}$ a code for correcting single $b$-burst-deletion. Let $\phi(\cdot)$ be the function defined in Lemma II.2. Then by Lemma II.2, our goal boils down to recovering the values of $\phi\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $A(\mathbf{x})_{1}$ is $d$-regular, we have $\left|I_{j}^{\prime}\right|\leq db\log(n/b)$ for all $0\leq j\leq r(U(\mathbf{x})_{1})$. Therefore, each $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in $[0,N-1]$, where $N=2^{\log\log(n)+O_{q,d,b}(1)}$. Here, $O_{q,d,b}(1)$ denotes a constant dependent only on $q,d$ and $b$.
Define $\overline{\phi}(\mathbf{x})=\phi\left(\mathbf{x}_{I_{0}^{\prime}}\right)\phi\left(\mathbf{x}_{I_{1}^{\prime}}\right)\cdots\phi\left(\mathbf{x}_{I_{r(U(\mathbf{x})_{1})}^{\prime}}\right)\in[0,N-1]^{r(U(\mathbf{x})_{1})+1}$. According to (13), when given $\mathbf{x}^{\prime}$, the value of $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, sequence $\overline{\phi}(\mathbf{x})$ suffered two bursts (of length two) of erasures. To correct errors in $\overline{\phi}(\mathbf{x})$, let $\varphi(\cdot)$ be the function defined in Lemma II.7 and define
@@
Recall that we have $r(U(\mathbf{x})_{1})\leq n/b$. Then By Lemma II.7, $f(\mathbf{x})$ can be viewed as an integer in $[0,N_{1}-1]$, where $N_{1}=2^{\log n+4\log\log n+O_{q,d,b}(1)}$.
The following lemma follows from the above analysis, Lemmas II.2 and II.7.
@@
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f(\mathbf{x})$, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
@@
Recall that case (2) in Lemma IV.5 corresponds to case (2) in Lemma IV.3. Therefore, for case (2), instead of looking at $\mathbf{x}$ and $\mathbf{x}^{\prime}$, we look at $A(\mathbf{x})$ and $A(\mathbf{x}^{\prime})$. In this case, we have already decoded $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ by using $\eta\left(U(\mathbf{x})_{1}\right)$. Furthermore, by case (2) in Lemma IV.3, we found an interval $[c_{1},c_{2}]\subseteq[n/b]$ of length at most $3db\log(n/b)$, which contains the positions of the two deletions in $U(\mathbf{x})_{1}$ (and hence in $A(\mathbf{x})_{1}$). By II.2, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $[c_{1}-1,c_{2}]$, for all $2\leq i\leq b$. Therefore, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin[c_{1}-1,c_{2}]$.
Note that $\left|[c_{1}-1,c_{2}]\right|$=$\left|J\right|+1\leq 3db\log(n/b)+1$. Let $P=3db\log(n/b)+1$ and define
@@
Since $c_{1}$ and $c_{2}$ are known, we can find some $j_{0}$ such that $[c_{1}-1,c_{2}]\subseteq J_{j_{0}}$. Suppose $J_{j_{0}}=[d_{1},d_{2}]$. According to above discussion, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin J_{j_{0}}$. To decode $A(\mathbf{x})_{i}$ from $A(\mathbf{x}^{\prime})_{i}$, it remains to recover $A(\mathbf{x})_{i,J_{j_{0}}}$.It is easy to see that $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i,J_{j_{0}}}\right)$. Therefore, we can apply Lemma II.4 to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$.
Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. For each $11\leq j\leq\left\lceil n/bP\right\rceil-$ 1, define
$g_{j}\left(\mathbf{x}\right)=\begin{cases}\left(\xi_{1}\left(A(\mathbf{x})_{1,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q>2;\\ \left(\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q=2.\end{cases}$
@@
Since $\left|J_{j}\right|\leq 2P=6db\log\left(n/b\right)+2$ for all $j$, by Lemma II.4, each $\xi_{1}\left(A(\mathbf{x})_{i,J_{j}}\right)$ is a binary vector of length $7\left\lceil\log q\right\rceil\log\log n$ $+O_{q,d,b}(1)$. Therefore, each $g_{j}(\mathbf{x})$ is a binary vector of length at most $7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)$ (when $q>2$) or $7(b-1)\log\log n+O_{d,b}(1)$ (when $q=2$). Therefore, we can view $g_{j}(\mathbf{x})$ as an integer in $\left[0,N_{2}-1\right]$, where
$N_{2}=\begin{cases}2^{7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)},\mbox{ if }q>2,\\ 2^{7(b-1)\log\log n+O_{d,b}(1)},\mbox{ if }q=2.\\ \end{cases}$ (16)
@@
For case (2) in Lemma IV.5 (and hence case (2) in Lemma IV.3), given $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$, we can uniquely recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $\mathbf{x}^{\prime}$, for each $2\leq i\leq b$. In particular, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
@@
Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O_{q,b}(1)$:
@@
Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma IV.5, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $13bd\log(n/b)+b-$ 1, such that $D_{1},D_{2}\subseteq J$. For the former case, Lemma IV.6 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. For the latter case, Lemma IV.7 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. Therefore, $\mathcal{C}_{1}$ can correct two $b$-burst-deletions.
By Lemma IV.4, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $A(\mathbf{x})_{1}$ is $d$-regular. It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, and $c_{1}$, such that

### v5

@@ 2. $(2)$
    or find an interval $J\subseteq[n]$ of length at most $3bd\log(n/b)+b-1$, such that $D_{1},D_{2}\subseteq J$.
@@
Since $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, we have $u(\mathbf{x}^{\prime})=u(\mathbf{x})_{[n]\setminus(D_{1}\cup D_{2})}$. It follows that $U(\mathbf{x}^{\prime})_{1}\in\mathcal{B}_{2}\left(U(\mathbf{x})_{1}\right)$. Suppose that $U(\mathbf{x}^{\prime})_{1}$ is obtained from $U(\mathbf{x})_{1}$ by deleting $U(\mathbf{x})_{1,k_{1}}$ and $U(\mathbf{x})_{1,k_{2}}$. Recall that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. This implies that $U(\mathbf{x}^{\prime})_{1}$ is obtained from $u(\mathbf{x})_{I_{n,b}}$ by deleting $u(\mathbf{x})_{1+(k_{1}-1)b}$ and $u(\mathbf{x})_{1+(k_{2}-1)b}$.
Since $U(\mathbf{x})_{1}$ is $d$-regular, according to Lemma IV.2, we can recover $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ with the help of $\eta\left(U(\mathbf{x})_{1}\right)$. Then we can know the number of runs in $U(\mathbf{x})_{1}$ and $I_{j}$ for all $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$. In addition, by Lemma IV.3, we can
@@
  • or find an interval $c,d]\subseteq[n/b]$ of length at most $3d\log(n/b)$ such that $k_{1},k_{2}\in[c,d]$.
Similar to the discussion in the paragraph above Lemma IV.5, for the first case, we have $D_{1}\subseteq\left[p_{j_{1}-1}b-b+2,p_{j_{1}}b\right]=I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}}b\right]=I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. For the second case, we have $D_{1},D_{2}\subseteq J\triangleq[(c-2)b+2,db]\cap[n]$. Now the proof is completed. ∎
From now on, we assume that $\mathbf{x}\in\Sigma_{q}^{n}$ and $A(\mathbf{x})_{1}$ is $d$-regular, where $d$ will be specified later. Let $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$, where $D_{1},D_{2}\subseteq[n]$ are two unknown disjoint intervals of length $b$. Suppose that $\eta\left(U(\mathbf{x})_{1}\right)$ is given. The first step of decoding $\mathbf{x}$ from $\mathbf{x}^{\prime}$ is to decode $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$. After this step, we can know the number of runs in $U(\mathbf{x})_{1}$ and $I_{j}$ for all $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$. Then by Lemma IV.5, one can approximately locate the two bursts in $\mathbf{x}$ if $\eta\left(U(\mathbf{x})_{1}\right)$ is given. Based on the two cases in Lemma IV.5, we split our discussion into Section IV-B and Section IV-C. Our codes for correcting two $b$-burst-deletions are presented in Theorem IV.1, whose proof is divided into Lemmas IV.6 and IV.7. In Section IV-B and Section IV-C, we follow notations defined in this subsection.
@@
Let $m>b$. Suppose $\mathbf{u}\in\Sigma_{q}^{m}$ and $\mathbf{v}$ is obtained from $\mathbf{u}$ by single $b$-burst-deletion. Then for any $b\leq i\leq m-b$, we have $\mathbf{v}_{[1,i-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[1,i]}\right)$ and $\mathbf{v}_{[i+1,m-b]}\in\mathcal{B}_{1}^{b}\left(\mathbf{u}_{[i+1,m]}\right)$.
@@
In particular, $\mathbf{x}_{I_{j}^{\prime}}$ is known to us for any $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, to decode $\mathbf{x}$ from $\mathbf{x}^{\prime}$, it is sufficient to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
@@
Since $j_{2}-j_{1}\geq 2$, deletions in $\mathbf{x}_{I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}}$ do not affect $\mathbf{x}_{I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}}$. As a result, we have $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}\right)$ (i.e., $\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$ is obtained from $\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}$ by single $b$-burst-deletion) and $\mathbf{x}^{\prime}_{\left[(p_{j_{2}-2}-1)b+1,p_{j_{2}}b-2b\right]}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}}\right)$ due to the fact that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}\subseteq I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}\subseteq I_{j_{2}-1}^{\prime}\cup I_{j_{2}}^{\prime}$. Applying IV.1 to $\mathbf{u}=\mathbf{x}_{I_{j_{1}-1}^{\prime}\cup I_{j_{1}}^{\prime}}$ (i.e., the concatenation of $\mathbf{x}_{I_{j_{1}-1}}$ and $\mathbf{x}_{I_{j_{1}}}$), $\mathbf{v}=\mathbf{x}^{\prime}_{\left[p_{j_{1}-2}b+1,p_{j_{1}}b-b\right]}$, $m=\left(p_{j_{1}}-p_{j_{1}-2}\right)b$ and $i=\left(p_{j_{1}-1}-p_{j_{1}-2}\right)b$, we conclude that
@@
Recall that the range of $j_{1}$ is $\left[1,r(U(\mathbf{x})_{1})\right]$. To make (14) valid when $j_{1}=1$, we let $\mathbf{x}_{I_{0}^{\prime}}=0^{b}$ (the sequence consisting of $b$ symbols 0) and $\mathbf{x}^{\prime}_{\left[p_{-1}b+1,-b\right]}$ be an empty sequence.
By (14) and (15), to recover $\mathbf{x}_{I_{j}^{\prime}}$ for $j\in\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$, we need to apply to each $\mathbf{x}_{I_{j}^{\prime}}$ a code for correcting single $b$-burst-deletion. Let $\phi(\cdot)$ be the function defined in Lemma II.2. Then by Lemma II.2, our goal boils down to recovering the values of $\phi\left(\mathbf{x}_{I_{j_{1}-1}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{1}}^{\prime}}\right)$, $\phi\left(\mathbf{x}_{I_{j_{2}-1}^{\prime}}\right)$ and $\phi\left(\mathbf{x}_{I_{j_{2}}^{\prime}}\right)$. Since $U(\mathbf{x})_{1}$ is $d$-regular and each $U(\mathbf{x})_{1,I_{j}}$ is a run of $U(\mathbf{x})_{1}$, we have $\left|I_{j}\right|\leq d\log(n/b)$ and hence $\left|I_{j}^{\prime}\right|\leq db\log(n/b)$ for all $0\leq j\leq r(U(\mathbf{x})_{1})$. Therefore, each $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ can be viewed as an integer in [0,N1][0,N-1], where $N=2^{\log\log(n)+O_{q,d,b}(1)}$. Here, $O_{q,d,b}(1)$ denotes a constant dependent only on $q,d$ and $b$.
Define $\overline{\phi}(\mathbf{x})=\phi\left(\mathbf{x}_{I_{0}^{\prime}}\right)\phi\left(\mathbf{x}_{I_{1}^{\prime}}\right)\cdots\phi\left(\mathbf{x}_{I_{r(U(\mathbf{x})_{1})}^{\prime}}\right)\in[0,N-1]^{r(U(\mathbf{x})_{1})+1}$. According to (13), when given $\mathbf{x}^{\prime}$, the value of $\phi\left(\mathbf{x}_{I_{j}^{\prime}}\right)$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$. Therefore, sequence $\overline{\phi}(\mathbf{x})$ suffered two bursts (of length two) of erasures. To correct errors in $\overline{\phi}(\mathbf{x})$, let $\varphi(\cdot)$ be the function defined in Lemma II.6 and define
@@
Recall that we have $r(U(\mathbf{x})_{1})\leq n/b$. Then By Lemma II.6, $f(\mathbf{x})$ can be viewed as an integer in $0,N_{1}-1]$, where $N_{1}=2^{\log n+4\log\log n+O_{q,d,b}(1)}$.
The following lemma follows from the above analysis, Lemmas II.2 and II.6.
@@
Suppose that we have found $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, where $j_{2}-j_{1}\geq 2$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq I_{j_{2}}^{L}\cup I_{j_{2}}^{\prime}$. Given $f(\mathbf{x})$, we can efficiently recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
@@
Recall that case (2) in Lemma IV.5 corresponds to case (2) in Lemma IV.3. Therefore, for case (2), instead of looking at $\mathbf{x}$ and $\mathbf{x}^{\prime}$, we look at $A(\mathbf{x})$ and $A(\mathbf{x}^{\prime})$. In this case, we have already decoded $U(\mathbf{x})_{1}$ from $U(\mathbf{x}^{\prime})_{1}$ by using $\eta\left(U(\mathbf{x})_{1}\right)$. Furthermore, by case (2) in Lemma IV.3, we found an interval $c_{1},c_{2}]\subseteq[n/b]$ of length at most $3d\log(n/b)$, which contains the positions of the two deletions in $U(\mathbf{x})_{1}$ (and hence in $A(\mathbf{x})_{1}$). By II.2, the two deletions in $A(\mathbf{x})_{i}$ occurred in the interval $c_{1}-1,c_{2}]$, for all $2\leq i\leq b$. Let $A(\mathbf{x})_{i,k}$ be the $k$-th coordinate in the $A(\mathbf{x})_{i}$. Then $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin[c_{1}-1,c_{2}]$. In fact, it holds that $A(\mathbf{x})_{i,k}=A(\mathbf{x}^{\prime})_{i,k}$ when $1\leq k<c_{1}-1$ and $A(\mathbf{x})_{i,k}=A(\mathbf{x}^{\prime})_{i,k-2}$ when $c_{2}<k\leq n/b$.
Let $P=3d\log(n/b)+1$. Define
@@
Since $c_{1}$ and $c_{2}$ are known and $\left|[c_{1}-1,c_{2}]\right|\leq 3d\log(n/b)+1$, we can find some $j_{0}$ such that $c_{1}-1,c_{2}]\subseteq J_{j_{0}}$. Suppose $J_{j_{0}}=[d_{1},d_{2}]$. According to above discussion, $A(\mathbf{x})_{i,k}$ is known to us for all $k\notin J_{j_{0}}$. To decode $A(\mathbf{x})_{i}$ from $A(\mathbf{x}^{\prime})_{i}$, it remains to recover $A(\mathbf{x})_{i,J_{j_{0}}}$. It is easy to see that $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$ is obtained from $A(\mathbf{x})_{i,J_{j_{0}}}$ by two deletions. Therefore, we can apply Lemma II.4 to recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $A(\mathbf{x}^{\prime})_{i,[d_{1},d_{2}-2]}$.
Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. For each $1\leq j\leq\left\lceil n/bP\right\rceil-1$, let function $g_{j}(\mathbf{x})$ be defined in (16).
$g_{j}\left(\mathbf{x}\right)=\begin{cases}\left(\xi_{1}\left(A(\mathbf{x})_{1,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q>2;\\ \left(\xi_{1}\left(A(\mathbf{x})_{2,J_{j}}\right),\xi_{1}\left(A(\mathbf{x})_{3,J_{j}}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b,J_{j}}\right)\right),\mbox{ if }q=2.\end{cases}$ (16)
@@
Since $\left|J_{j}\right|\leq 2P=6d\log\left(n/b\right)+2$ for all $j$, by Lemma II.4, each $\xi_{1}\left(A(\mathbf{x})_{i,J_{j}}\right)$ is a binary vector of length $7\left\lceil\log q\right\rceil\log\log n$ $O_{q,d,b}(1)$. Therefore, each $g_{j}(\mathbf{x})$ is a binary vector of length at most $7b\left\lceil\log q\right\rceil\log\log n+O_{q,d,b}(1)$ (when $q>2$) or $7(b-1)\log\log n+O_{d,b}(1)$ (when $q=2$). Therefore, we can view $g_{j}(\mathbf{x})$ as an integer in $\left[0,N_{2}-1\right]$, where
[…]
@@
For case (2) in Lemma IV.5 (and hence case (2) in Lemma IV.3), given $h^{(0)}(\mathbf{x})$ and $h^{(1)}(\mathbf{x})$, we can uniquely recover $A(\mathbf{x})_{i,J_{j_{0}}}$ from $\mathbf{x}^{\prime}$, for all $i$. In particular, we can recover $\mathbf{x}$ from $\mathbf{x}^{\prime}$.
@@
Then $\mathcal{C}_{1}$ is a 22-$b$-burst-deletion correcting code. Furthermore, when $(q,n,d)$ satisfies one of the following conditions, there is some $a$, $b$, $c_{0}$ and $c_{1}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O_{q,b}(1)$:
@@
Suppose $\mathbf{x}\in\mathcal{C}_{1}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Since $U(\mathbf{x})_{1}$ is $d$-regular, by Lemma IV.5, given $\eta\left(U(\mathbf{x})_{1}\right)$, we can either find $1\leq j_{1}<j_{2}\leq r\left(U(\mathbf{x})_{1}\right)$, such that $D_{1}\subseteq I_{j_{1}}^{L}\cup I_{j_{1}}^{\prime}$ and $D_{2}\subseteq\left[p_{j_{2}-1}b-b+2,p_{j_{2}-1}b\right]\cup I_{j_{2}}^{\prime}$; or find an interval $J\subseteq[n]$ of length at most $3bd\log(n/b)+b-1$, such that $D_{1},D_{2}\subseteq J$. For the former case, Lemma IV.6 assert that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. For the latter case, Lemma IV.7 asserts that $\mathbf{x}$ can be decoded from $\mathbf{x}^{\prime}$. Therefore, $\mathcal{C}_{1}$ can correct two $b$-burst-deletions.
By Lemma IV.4, there are at least $q^{n-1}$ sequences $\mathbf{x}\in\Sigma_{q}^{n}$ such that $A(\mathbf{x})_{1}$ is $d$-regular. It then follows from the pigeonhole principle that there are some $a$, $b$, $c_{0}$, and $c_{1}$, such that